May 28, 2009 Calorimetry

From GlueXWiki
Jump to: navigation, search

Teleconference Time: 11:00 a.m. EST

  • EVO session
  • Phone:
    • +1-800-377-8846 : US
    • +1-888-276-7715 : Canada
    • +1-302-709-8424 : International
    • then enter participant code: 39527048# (remember the "#")

Items for followup from previous meeting(s)

  1. Resolution studies (Blake)
    • Develop criteria for evaluating resolution studies
    • Study impact of inactive material at the end of BCAL on acceptance.
    • Check what configuration (thresholds, step sizes, stack? parameters for secondaries, etc) is being used for the Hall D simulations (Regina)
  2. Review options for import/export to Chile (loans, donations, transfers, etc.)
  3. Documentation on fiber first-article
  4. Design goals for SiPM electronics
  5. Preliminary specifications for Bcal electronics readout

Documents to Review

Tentative Agenda

  1. Announcements
  2. Review of Action Items
  3. Calibration
  4. FCAL Update
  5. Simulations
  6. BCAL Fabrication
  7. BCAL Readout
    • Update from USM (Will)
    • Readout specifications using SiPMs (Fernando)
    • Rates in the BCAL [1] (Sasha) and Beni
    • Light guide efficiency for FM PMTs (GlueX-doc-1279 pp 12-19) (Elton)
    • Review of readout options (tentative July 14) discussion


Attending (from memory, I might have missed someone) USM: Sergey Kuleshov, Will Brooks, Alam Toro, Pedro Toledo, Eduardo Mora, JLab: Elton, Fernando, Beni, Sasha, Eugene, UofR; Zisis, George, Andrei, IU: Ryan, Claire

  1. Announcements
    • Over the next month, candidates will start interviewing for Physics APM and Hall D group leader. First candidate interviews today.
  2. Review of Action Items
    1. Import/export issues in Chile. Issues are being clarified, but work in this direction will continue.
      • One large scientific instrument paid for by UN had tariffs waived
      • Taxes and tariffs are reduced or waived when the University pays for related research activities. However, many grants (such as Will's) do not go through the university accounting system.
      • Business and scientific personnel continue to investigate the nuisances.
      • Zisis: University is investigating how to ship the baby calorimeter to Chile. The 260 lbs 2 ft section arrived from Alberta this week. Investigating options for cutting it into smaller sections. Will suggested that the cost for shipping the entire segment be requested. It is possible that the shipping cost is not directly related to weight.
  3. Calibration
    • Christina has sent a 20 p. report on calibration to Zisis for feedback and should be placed on the portal soon.
  4. FCAL Update (Ryan)
    • Sixty-four element array under construction to resolve mechanical issues
    • Jake finalizing timing studies with Flash ADC
    • Claire is working on the calibration using pi0s, adding the rad hard glass and boundary between center and outer Fcal.
    • Prototype for slow controls under development, Elton suggested that they make sure that Elliott is aware of their work.
  5. Simulations (Andrei)
    • Using detailed MC to generate photon showers between 0.2 and 1 GeV at 90 deg. Will then work at 20 deg.
    • Elton suggested that they generate showers at the extremes (i.e. 0.04 and 2 GeV) to make sure that they cover the range of expected sizes.
    • Adding attenuation length in calorimeter as well as photostatistics.
    • Using nominal segmentation with 2x2 outer and 6x4 inner configuration.
  6. BCAL Fabrication
    1. Contract (George)
      • Contract response has been completed and reviewed with Helen Yun from their research office.
      • Should be ready for signature at UofR tomorrow for responding to JLab.
      • Some detailed quotes are still in the works, including ones for the second press frame.
    2. Preparations for construction (Zisis)
      • Received notice that six students are applied for coop positions for the fall, four are interested in year appointments. This is good news.
      • Tentative offers will be made in the next few weeks.
      • Room preparations continue. They are not slowing down operations since only one press frame is operational.
      • Received press tables and equipment from Alberta earlier this week.
      • All available lead hass been cut (keeping one out of eleven rolls as spare). This is be enough for about 2/3 of the construction prototype.
      • Tim: Order for lead should go out in the next few days, and first lead deliveries are expected in 30 days. Will coordinate with Zisis sizes that are most needed.
      • Will: Noted from CLAS experience both oxides and carbonates can accumulate. One type is worse than the other (will send references to UofR). His recollection is that CLAS used a HEPA vacuum cleaner and then cleaned the lead with alcohol and did not use chemical cleaners.
      • Swager has been installed and aligned.
      • Orders for epoxy and consumables has been put out.
      • No news of fiber deliveries (Elton will check)
      • Discussion/tests are ongoing relative to the use of vaseline or other methods.
      • Investigating possibilities of using soft PVC to store fibers. Will asked whether migration of plasticizers is a problem.
      • Elton will also check from Kuraray if they have any comments on where the fibers are stored.
    3. Report on First-article fibers (Zisis)
  7. BCAL Readout
    1. Update from USM (Will)
      • Alam: reported on getting a CAD system to establish location of light guides
      • Eduardo (Manuel) preparing report on characteristics of fibers to work with pmts and SiPMs
      • Pedro: working on FM pmts: reviewing documents from Athens and FIU to determine optimal location in the magnetic field . The maximum gain occurs at about 40 deg. (Elton suggests that setting them at zero degrees to the field will have the least sensitivity to field uncertainties.)
    2. Electrical specifications using SiPMs (Fernando)
      • Draft of electrical specifications is essentially complete and has been circulated for comments. Fernando will send this out to SensL and Hamamatsu this afternoon.
    3. Rates in the BCAL [2] (Sasha) and Beni
      • Sasha and Beni have computed the number of visible photons expected in the upstream and downstream pmts of the Bcal for events generated with PYTHIA. They conclude the expected dynamic range is about 1:200. This is a factor of 3 lower than what is presently used by Fernando, which gives some headroom.
    4. Light guide efficiency for FM PMTs (GlueX-doc-1279 pp 12-19) (Elton)
      • Elton has generated some reference configurations for light guides using a 90 deg bend. The collection efficiency is close to the geometrical ratio of areas, but is shown to be relatively unbiased relative to input position. He will complete some adjustments to the configurations and then make them available to USM for more detailed studies.
      • Will inquired about other possible configurations such as curved light guides and other orientations in the field. These can certainly be investigated as one considers the various pros and cons,
    5. Review of readout options (tentative July 14) discussion
      • It has been suggested that we have at least one outside reviewer and Elton will contact Stefano Miscetti to see if he is available for the review.